OpenStudy (anonymous):

x^2-x=3 and x^2+5=4x

OpenStudy (anonymous):

Are you trying to solve simultaneous equations?

OpenStudy (anonymous):

If you want find what values for X solve both equations then you are going to have to use the quadratic formula

OpenStudy (dumbcow):

set both equations equal to 0 x^2-x-3 =0 x^2 -4x+5=0 Now set them equal to each other to solve for x x^2-x-3 = x^2 -4x+5 combine like terms 3x = 8 x = 8/3